OpenStudy (anonymous):

help me solve |3/5w-2|+5=3

OpenStudy (gorv):

subtract 5 from both side

OpenStudy (gorv):

|3/5w-2|=3-5=-2 but modulus dont give negative value

OpenStudy (gorv):

no solution